A skateboarder travels on a horizontal surface with an initial velocity of 4.4 m/s toward the south and a constant acceleration of 1.7 m/s2 toward the east. Let the x direction be eastward and the y direction be northward, and let the skateboarder be at the origin at t=0 .

Part A

Part complete

What is her x position at t=0.70s ?

Express your answer to two significant figures and include appropriate units.

--

SubmitPrevious AnswersRequest Answer

Part B

Part complete

What is her y position at t=0.70s ?

Express your answer to two significant figures and include appropriate units.

--

nothing Value nothing Units

SubmitPrevious AnswersRequest Answer

Part C

Part complete

What is her x velocity component at t=0.70s ?

Express your answer to two significant figures and include appropriate units.

--

nothing Value nothing Units

SubmitPrevious AnswersRequest Answer

Part D

Part complete

What is her y velocity component at t=0.70s ?
Answer #1

a) X = Uxt + 0.5axt^2

X = 0 + 0.5*1.7*0.7^2

X = 0.42 m

b) Y = Uyt + 0.5at^2

Y = 4.4 *0.7 + 0

= 3.1 m

c) Vx = Ux + at

Vx = 0 + 1.7*0.7

Vx = 1.2 m/s ( east)

d) Vy = Uy +at

Vy = 4.4 + 0

Vy = 4.4 m/s (south)